OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the graphing method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (2 points) Part 3: Explain what the solution represents on the graph. (1 point) 3x + 6y = 18 2x + 5y = 20 I'm so lost :(

jimthompson5910 (jim_thompson5910):

plug in x = 0 to find y 3x + 6y = 18 3(0) + 6y = 18 0+6y = 18 6y = 18 y = 18/6 y = 3 so you have one point (0, 3) on the line 3x + 6y = 18 to find another point, plug in y = 0 and solve for x

OpenStudy (anonymous):

So if x=6, what do I do next?

jimthompson5910 (jim_thompson5910):

(6, 0) is another point on 3x + 6y = 18

jimthompson5910 (jim_thompson5910):

you then plot the two points (0, 3) and (6,0) then draw a straight line through them

OpenStudy (anonymous):

So how does that get me to the solution?

jimthompson5910 (jim_thompson5910):

you repeat the same steps to graph 2x + 5y = 20

jimthompson5910 (jim_thompson5910):

where the two lines cross is where your solution point is

OpenStudy (anonymous):

Ahh I see!

jimthompson5910 (jim_thompson5910):

that's great

OpenStudy (anonymous):

do you thinkyou could help me with the elimination method?

jimthompson5910 (jim_thompson5910):

which variable do you want to eliminate

OpenStudy (anonymous):

actually its a different problem, sorry my dog had to be let out

OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. (4 points) Part 2: Provide the solution to the system. (2 points) 5x – 3y = 4 3x – 5y = –4

jimthompson5910 (jim_thompson5910):

which variable do you want to eliminate

OpenStudy (anonymous):

I have no idea, which would probably be the simplest?

jimthompson5910 (jim_thompson5910):

either one is about the same amount of work, so just pick one

OpenStudy (anonymous):

the first

jimthompson5910 (jim_thompson5910):

ok if we just add 5x and 3x we get 8x what we want is the two to add to 0x so the x terms get eliminated

jimthompson5910 (jim_thompson5910):

how are we going to make the x terms add to 0x? well we could multiply the 5x by 3 to get 15x and we could multiply the 3x by -5 to get -15x notice how 15x + (-15x) = 15x - 15x = 0x = 0 making sense so far?